1. Find the Voltages at Nodes V1 and V2
2. Find Current through R1
3. Find The Power Delivered/Sunk by I1
4. Find the current with direction through R2
R2 W 90 15A R3 Dron 10A 30 3120
Answer #1

Given Circuit loop3 un ISA W & 122 & 4 loop! y loop2 We can solve given circuit by using loop From loop! I = 10A From loop2 39Iz- 974 = 135 1513 + ay = -30 24 I3 = 105 Iz= 105 2 4 375 Substituting Iz in €90 I4 15-73 - Iz-I4 = 15 ; -I4 = 15-4.375 Iy =2) Current through Ri it through R = 14:375 A 3) Power delivered or sunk by Is By obseoving the circuit power Power delivered